# BNU-HKBU UIC Zhuhai, China

### v2

@@ II Related Work
Recently, the most widely used GAN [1] based image fusion methods is to contrast two discriminators for fused images and both source images [2, 3]. For example, FusionGAN for IVIF task, in which the generator aims to generate fused images that major infrared intensities together with additional visible gradients and the discriminator aims to force the fused image to have more details existing just in the visible image [4, 5]. The adversarial game of it is designed between fused image and phase-contrast image [6]. AE-based methods focuses on the elaborated reconstruction and fusion loss functions [7, 8, 9, 10]. For instance, [7] first adopted an AE-based model in the IVIF task. Moreover, [8, 11] proposed for high/low-frequency decomposition strategy aims to extract features more exactly for IVIF task in recent years.
@@ IV-A1 Datasets and metrics
In the first phase, we trained 80 epochs for both the autoencoder and discriminator, while in the second phase, the model was trained with 140 epochs. We use the Adam optimizer for the autoencoder and the RMSProp optimizer for the discriminator with an initial learning rate of 1e-4 and decreasing by 0.5 every 20 epochs. The training samples are randomly cropped into 128 × 128 patches in preprocessing stage and the batch size is set to 16. In the MMOD downstream testing, the generated 4200 fused images is partitioned into training, validation, and test sets, with an 8:1:1 ratio for a YOLOv8n [26], which is recognized as the SOTA detector. The number of epochs is set to 320, batch size is set to 8, the optimizer is performed in SGD, and the initial learning rate is set to 1e-2, respectively.
@@ VI Acknowledgement
Our work was supported in part by the Guangdong Provincial Key Laboratory of Interdisciplinary Research and Application for Data Science, BNU-HKBU United International College, project code 2022B1212010006, and in part by Guangdong Higher Education Upgrading Plan (2021-2025) of ”Rushing to the Top, Making Up Shortcomings and Strengthening Special Features” with UIC research grant UICR0400006-24.

### v3

# Guangdong Provincial Key Lab of Interdisciplinary Research and Application for Data Science

@@ II Related Work
Recently, the most widely used GAN [1] based image fusion methods is to contrast two discriminators for fused images and both source images [2, 3]. For example, FusionGAN for IVIF task, in which the generator aims to generate fused images that major infrared intensities together with additional visible gradients and the discriminator aims to force the fused image to have more details existing just in the visible image [4]. The adversarial game of it is designed between fused image and phase-contrast image [5]. AE-based methods focuses on the elaborated reconstruction and fusion loss functions [6, 7, 8]. For instance, [6] first adopted an AE-based model in the IVIF task. Moreover, [7, 9] proposed for high/low-frequency decomposition strategy aims to extract features more exactly for IVIF task in recent years.
@@ IV-A1 Datasets and metrics
In the first phase, we trained 80 epochs for both the autoencoder and discriminator, while in the second phase, the model was trained with 140 epochs. We use the Adam optimizer for the autoencoder and the RMSProp optimizer for the discriminator with an initial learning rate of 1e-4 and decreasing by 0.5 every 20 epochs. The training samples are randomly cropped into 128 × 128 patches in preprocessing stage and the batch size is set to 16. In the MMOD downstream testing, the generated 4200 fused images is partitioned into training, validation, and test sets, with an 8:1:1 ratio for a YOLOv8n [24].
@@ VI Acknowledgement
Our work was supported in part by the Guangdong Provincial Key Laboratory of IRADS (2022B1212010006) and in part by Guangdong Higher Education Upgrading Plan (2021-2025) with No.of UICR0400006-24, UICR0400005-24.